OpenStudy (anonymous):

you have 288 feet of fencing to enclose a rectangular plot that borders on a river. if you do not fence the side along the river, what is the largest area that can be enclosed?

OpenStudy (anonymous):

it needs to look like a square

OpenStudy (anonymous):

or as close as a square as it can get

OpenStudy (anonymous):

so 96 on the 3 sides

OpenStudy (anonymous):

first the answer take half of the fence and put it parallel to the river other remaining half gets split between the two sides

OpenStudy (anonymous):

@haha1231233123 unfortunately that is not right

OpenStudy (anonymous):

and the area of the square is 96 times 96 = 9216.sqft

OpenStudy (anonymous):

wouldn't you use a = 2w + l ?

OpenStudy (anonymous):

144 by 72

OpenStudy (anonymous):

you have \[2x+y=288\] so \(y=288-2x\) and \[A=xy\] so \[A(x)=x(288-2x)=288x-2x^2\]

OpenStudy (anonymous):

max is at the vertex, and first coordinate of the vertex is \(-\frac{b}{2a}=\frac{288}{4}=72\) as promised